OpenStudy (anonymous):

2.Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (cwrw238):

he can check his answer by plugging in x=1 into the original eqauation to see if it fits

OpenStudy (cwrw238):

if you check it then LHS = 3 * 9 = 27 and RHS = 4*6 = 24 so x=1 is not correct

OpenStudy (anonymous):

thanks, but where is the mistake

OpenStudy (cwrw238):

there is an error in the second line - can you spot it?

OpenStudy (cwrw238):

if you distribute 4 over (7x - 1) what do you get?

OpenStudy (anonymous):

you get 28x - 4

OpenStudy (cwrw238):

right - theres the mistake

OpenStudy (anonymous):

ok now I see it

OpenStudy (cwrw238):

so we have 9x+18 = 28x−4 9x - 9x + 18 = 28x - 9x - 4 18 = 19x - 4

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

which equals no solution, thx for all your help

OpenStudy (cwrw238):

no - its not no solution - there is a solution

OpenStudy (cwrw238):

18 = 19x - 4 18 + 4 = 19x -4 + 4 22 = 19x 22/19 = x

OpenStudy (anonymous):

ohhhhhhhhhh ok thx

OpenStudy (cwrw238):

yw